Aaron, Betsy,and Charita work part time at the public library. Betsy works 4hours more each week than Aaron, and together they work as many as Charita. How long does each person work if their total is 45h?
Found 2 solutions by funmath, venugopalramana:
Answer by funmath(2933)

Let the amount that Aaron works=x
Betsy works 4 hrs more than Aaorn=x+4
Charisa works as much as they both do put together=(x)+(x+4)=2x+4
They work a total of =45
Problem:
(x)+(x+4)+(2x+4)=45
(1+1+2)x+(4+4)=45
4x+8=45
4x+8-8=45-8
4x=37
4x/4=37/4
x=9.25 hrs
Aaron works:x=9.25 hrs, or 9hrs and 15 min
Betsy works:x+4=9.25+4=13.25 hrs, or 13 hrs and 15 minutes
Charisa:2x+4=2(9.25)+4=18.5+4=22.5 hrs, or 22 hrs and 30 minutes
Happy Calculating!!!!!

Answer by venugopalramana(3286)

Aaron ( A SAY), Betsy ( B SAY ),and Charita ( C SAY )work part time at the public library. Betsy works 4hours more each week than Aaron
B = A + 4 .................................1
, and together they work as many as Charita.
A + B = C....................2..........SUBSTITUTING FOR B FROM EQN.1 IN EQN.2
A+A+4=C
C= 2A+4...........................................3
How long does each person work if their total is 45h?
A+B+C = 45.........................4....SUBSTITUTING FOR B AND C FROM EQNS.1 AND 3
A+A+4+2A+4=45
4A=45-8=37
A=37/4 = 9.25
B = 9.25+4 = 13.25
C = 2*9.25 + 4 =22.5